OpenStudy (anonymous):

i need some help with algebra 2 medal for best answer.

OpenStudy (anonymous):

I'm in algebra 1, but i'll do what i can.

OpenStudy (anonymous):

im in connexus and need lesson 8 unit 1

OpenStudy (anonymous):

Can u show the prob?

OpenStudy (anonymous):

find any points of discontinuity for the rational function. y=x-8/x^2+6x-7

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

@SiennaSurfer

OpenStudy (amoodarya):

whats the question ?

OpenStudy (anonymous):

find any points of discontinuity for the rational function. y=x-8/x^2+6x-7

OpenStudy (amoodarya):

find the roots of denominator

OpenStudy (anonymous):

idk how to do that...

OpenStudy (anonymous):

:-( hello?

OpenStudy (amoodarya):

ok my net was in problem now lets find roots x^2+6x-7 =0 (x+7)(x-1)=0 because -1+7=6 -1 *6=-7 can you go on ?

OpenStudy (amoodarya):

x-8 y=---------- (x+7)(x-1)

OpenStudy (amoodarya):

so x=-7 , +1 points of discontinuity

OpenStudy (anonymous):

so the 7 is negative?

OpenStudy (amoodarya):

(x+7)(x-1) =x^2 +(7-1)x +(-1)(7) =x^2+6x -7